Sonia invests some of her money at 5% and twice as much at 5.5%. If her total income is $125, find the amount invested at each rate.
Answer by mananth(16946)

-------------- 5.00% per annum ------------- x
------------- 5.50% per annum------------y
y=2x
Interest----- $125.00

5.00% x + 5.50% * 2 x = $125.00
0.05 x + 0.06 * 2 x = $125.00
0.05 x + 0.11 x = $125.00
0.16 x = $125.00
/ 0.16
x= $781.25 ----------------- @5%
2 x = $1,562.50------------- @5.5%

CHECK
$39.06 + $85.94 = $125.00